OpenStudy (anonymous):

The half-life of a certain radioactive material is 42 days. An initial amount of the material has a mass of 49 kg. Write an exponential function that models the decay of this material. Find how much radioactive material remains after 8 days. Round your answer to the nearest thousandth.

OpenStudy (anonymous):

I know this has something to do with this formula I just don't know how to apply it? \[y = ab^{x} + k\]

OpenStudy (anonymous):

@Luigi0210

OpenStudy (anonymous):

The half-life of a certain radioactive material is 42 days. An initial amount of the material has a mass of 49 kg. Write an exponential function that models the decay of this material. : A = Ao*2^(-t/h), where: A = amt after t time Ao = initial amt (t=0) t = time of decay h = half life of substance : A = 49*2^(-t/42) : : Find how much radioactive material remains after 8 days. A = 49*2^(-8/42) A = 49*.8763 A = 42.9395 ~ 42.940, to the nearest thousandth.

OpenStudy (tkhunny):

The constant term, "k", has no business in this formulation. It neither grows nor deteriorates. No good.

OpenStudy (anonymous):

@jennaknight_03 Could you clarify on how you got all that? A = 49*2^(-t/42) How did you times 49 by 2? Where did you get the 2?

OpenStudy (anonymous):

im not sure about that 1. Ask @Luigi0210 about it

OpenStudy (tkhunny):

The half-life of a certain radioactive material is 42 days. An initial amount of the material has a mass of 49 kg. Write an exponential function that models the decay of this material. Find how much radioactive material remains after 8 days. Round your answer to the nearest thousandth. Given an initial amount, A0, we have: \((1/2)A_{0} = A_{0}e^{k\cdot 42}\). Solving for k, gives: \((1/2) = e^{k\cdot 42}\) \(-\ln(2) = k\cdot 42\) \(k = -\dfrac{\ln(2)}{42}\) Now, let's test it. Given \(A_{0} = 49\), we should have \(A_{42} = 49/2 = 24.5\) \(49e^{\left(-\dfrac{\ln(2)}{42}\right)\cdot 42} = 49e^{-\ln(2)} = 49\cdot\dfrac{1}{2} = 24.5\) Now, let's try it for t = 8. \(49e^{\left(-\dfrac{\ln(2)}{42}\right)\cdot 8} =\;\;\)??